04: The switch in the has been open a long time before closing at t=0. Find 1,(() for t> 0. 131 -62.5H 2.5 1F

Help me know please , I have an exam its time 20 minutes
Answer #1

t-o u 262.54 15V It to- opposite direction of suutch, - se C-0.c. + Vrol Isv 180-1 EL (0) = -Is - GMA 2.5k Vco) = ov. 910J =0.016 0.4x100 +2.5x1 3 9 V + + -15 2.5KXS { S 2:5*7*$? » asarid to uns toobe} -SME + Ski 2.54183 +0.4x10 st 0.016 s? + 1605с IS A S + alu f B s +80 (8.48012 A=-3x103 3 B = 18x 10 -1.2. 3 I (s) = - 3x10 18 X10 (1.2) (8+80)2 S stoo Using Inverse Lap

Please Rate the Answer